# A compatible finite element discretisation for moist shallow water equations

### v1

@@ 5.2 Moist Flow over an Isolated Mountain
Figure 2: Potential vorticity (defined by ×$\frac{\nabla\times(\bm{u}+f)}{D}$) at Day 50 in the flow over a mountain test. (a) the moist convective formulation, (b) the moist convective thermal formulation, (c) the moist thermal formulation, and (d) the moist convective pseudo-thermal formulation.
@@ 5.3 Moist Unstable Jet
There is no prognostic buoyancy variable in the moist convective formulation of this test so the initial velocity and depth in that formulation are the same as in the original jet in Galewsky et al. [26]. The $\theta$ that appears in the saturation function (39) is given by $\theta=-\frac{\Delta b\cos\phi}{g}$ (coming from the relationship between $b$ and $\theta$, with $\theta$ as given for the thermal versions of the test and $\Delta b$ again a constant $1\text{m s}^{-2}$). The initial vapour is, as before, $(1-\xi)q_{sat}$, with $\xi=0.02$.
Like the flow over a mountain test there is no analytical solution for this test, but rather the results are assessed on the basis of how realistically cloud and rain are produced. The development and evolution of the jet instability in all four frameworks is consistent with the dry simulation results in Galewsky et al. [26], and with the moist thermal results in [45]. In all four formulations we see cloud formation about 4.5 days into the simulation as cloud is produced in the vortices. By about 5.25 days the conversion from cloud to rain has begun, and by day 6 there is a clear pattern of cloud and rain matching the roll-ups of the jet.
Figure 4: Potential vorticity at Day 6 in the moist unstable jet test. (a) the moist convective formulation, (b) the moist convective thermal formulation, (c) the moist thermal formulation, and (d) the pseudo-moist convective thermal formulation.
@@
Figure 5: Accumulated rainfall at Day 6 in moist unstable jet test. (a) the moist convective formulation, (b) the moist convective thermal formulation, (c) the moist thermal formulation, and (d) the pseudo-moist convective thermal formulation. As in Figure 3, we plot $\text{log}_{10}(q_{r})$ rather than $q_{r}$ directly so that we can plot all four configurations on the same scale despite variation in rain quantities, and omit rain values of less than $1\times 10^{-6}$ from the plot.
@@ 6 Conclusions
The moist shallow water model provides a useful tool with which to investigate questions about physics-dynamics coupling in time-stepping schemes. Our compatible finite element discretisation of the general moist shallow water model offers a framework with which four different moist shallow water models can be interrogated and compared, including a new, previously unexplored moist convective pseudo-thermal shallow water model. We intend to follow this work up by using the moist shallow water equations to investigate how time steppers compare at coupling different processes in time. In addition to this, the test set-ups and reference solutions we have detailed here can be useful as validation and benchmarks for others working with moist shallow water models, irrespective of the numerical schemes employed.
@@
The research data supporting this publication are provided within this paper.
@@ Appendix A Physics schemes
Though all our tests are run using the three-state physics scheme from Zerroukat and Allen [7], we also implemented the one-way physics scheme described by Bouchut et al. [4]. This is the scheme previously used with the moist convective and the moist convective thermal shallow water equations. To compare the two schemes we ran the moist convective version of the flow over a mountain test using each scheme in turn, defining the same depth-dependent saturation function in each case. In the one-way physics scheme this saturation function is the threshold for the direct conversion from vapour to rain and is not reversible. The proportion above saturation converted, $\gamma_{r}$, is set at $1\times 10^{-3}$. In the three-state scheme the same saturation function dictates conversion between vapour and cloud, which is a reversible conversion. The proportion converted, $\gamma_{v}$, depends on $D$ and is given by Equation (52). Rain is produced when cloud exceeds a fixed threshold, $q_{precip}$, and the proportion converted, $\gamma_{r}$, is $1\times 10^{-3}$.
@@ Appendix B Choice of $\beta_{1}$ parameter
Figure 8: Comparison of the vorticity after 50 days in the moist convective thermal version of the flow over a mountain test using four different values for the $\beta_{1}$ parameter. (a) $\beta_{1}=1.6\text{m}$, (b) $\beta_{1}=1600\text{m}$, (c) $\beta_{1}=8500\text{m}$, and (d) $\beta_{1}=10000\text{m}$. $\beta_{2}$ is fixed throughout at $10g\text{ m s}^{2}$.

### v2

@@ 5.2 Moist Flow over an Isolated Mountain
Figure 2: Potential vorticity (defined by (×$\frac{(\nabla\times\bm{u})\cdot\bm{\hat{k}}+f}{D}$) at Day 50 in the flow over a mountain test. (a) the moist convective formulation, (b) the moist convective thermal formulation, (c) the moist thermal formulation, and (d) the moist convective pseudo-thermal formulation.
@@ 5.3 Moist Unstable Jet
There is no prognostic buoyancy variable in the moist convective formulation of this test so the initial velocity and depth in that formulation are the same as in the original jet in Galewsky et al. [26]. The $\theta$ that appears in the saturation function (35) is given by $\theta=-\frac{\Delta b\cos\phi}{g}$ (coming from the relationship between $b$ and $\theta$, with $\theta$ as given for the thermal versions of the test and $\Delta b$ again a constant 1 $1\text{ m s}^{-2}$). The initial vapour is, as before, $(1-\xi)q_{sat}$, with $\xi=0.02$.
Like the flow over a mountain test there is no analytical solution for this test, but rather the results are assessed on the basis of the appearance of expected physical features, such as rain being generated along the fronts as the jet rolls up. The development and evolution of the jet instability in all four frameworks is consistent with the dry simulation results in Galewsky et al. [26], and with the moist thermal results in [45]. In all four formulations we see cloud formation about 4.5 days into the simulation as cloud is produced in the vortices. By about 5.25 days the conversion from cloud to rain has begun, and by day 6 there is a clear pattern of cloud and rain matching the roll-ups of the jet.
Figure 4: Potential vorticity (defined by (×$\frac{(\nabla\times\bm{u})\cdot\bm{\hat{k}}+f}{D}$) at Day 6 in the moist unstable jet test. (a) the moist convective formulation, (b) the moist convective thermal formulation, (c) the moist thermal formulation, and (d) the pseudo-moist convective thermal formulation.
@@
Figure 5: Accumulated rainfall at Day 6 in the moist unstable jet test. (a) the moist convective formulation, (b) the moist convective thermal formulation, (c) the moist thermal formulation, and (d) the pseudo-moist convective thermal formulation. As in Figure 3, we plot $\text{log}_{10}(q_{r})$ rather than $q_{r}$ directly so that we can plot all four configurations on the same scale despite variation in rain quantities, and omit rain values of less than $1\times 10^{-6}$ from the plot.
@@ 6 Conclusions
The moist shallow water model provides a useful tool with which to investigate questions about physics-dynamics coupling in time-stepping schemes. Our compatible finite element discretisation of the general moist shallow water model offers a framework within which four different moist shallow water models can be interrogated and compared, including a new, previously unexplored moist convective pseudo-thermal shallow water model. We intend to follow this work up by using the moist shallow water equations to investigate how time steppers compare at coupling different processes in time. In addition to this, the test set-ups and example solutions we have detailed here can be useful as solutions to verify against for others working with moist shallow water models, irrespective of the numerical schemes employed.
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Physics schemes
Though all our tests are run using the three-state physics scheme from Zerroukat and Allen [7], we also implemented the one-way physics scheme described by Bouchut et al. [4]. This is the scheme previously used with the moist convective and the moist convective thermal shallow water equations. To compare the two schemes we ran the moist convective version of the flow over a mountain test using each scheme in turn, defining the same depth-dependent saturation function in each case. In the one-way physics scheme this saturation function is the threshold for the direct conversion from vapour to rain and is not reversible. The proportion above saturation converted, $\gamma_{r}$, is set at $1\times 10^{-3}\text{ s}^{-1}$. In the three-state scheme the same saturation function dictates conversion between vapour and cloud, which is a reversible conversion. The proportion converted, $\gamma_{v}$, depends on $D$ and is given by Equation (44). Rain is produced when cloud exceeds a fixed threshold, $q_{precip}$, and the proportion converted, $\gamma_{r}$, is $1\times 10^{-3}\text{ s}^{-1}$.
@@ Appendix B Choice of $\beta_{1}$ parameter
Figure 8: Comparison of the vorticity after 50 days in the moist convective thermal version of the flow over a mountain test using four different values for the $\beta_{1}$ parameter. (a) $\beta_{1}=1.6\text{ m}$, (b) $\beta_{1}=1600\text{ m}$, (c) $\beta_{1}=8500\text{ m}$, and (d) $\beta_{1}=10000\text{ m}$. $\beta_{2}$ is fixed throughout at $10g\text{ m s}^{-2}$.